OpenStudy (anonymous):

a gardener is planning to make a rectungular garden with an area of 80ft square. she has 12 yards of fencing to put ground the perimeter of the garden. what should the dimension of the garden be?

OpenStudy (anonymous):

uii.. hahah nakuha ko na ung ans.. :p

OpenStudy (anonymous):

let the to sides be x nd y, so arey=x*y=80 and x+y=12/2=6 yards, now convert it into feets= 18 feets so x=18-y puttting in 1st equation, y(18-y)=80 clearly the asnwer is 8 nd 10